OpenStudy (anonymous):

Molly finds two sticks and wants to build a kite for her little brother. The sticks are 16 inches and 22 inches long, and she plans on using them as the diagonals of the kite. Molly will need a minimum of _____ square inches of fabric to make the kite.

OpenStudy (dumbcow):

Area = 16*22/2 = 8*22 = 176

OpenStudy (anonymous):

thank u so much

OpenStudy (anonymous):

was it right?